# WaveDiffusion: Exploring Full Waveform Inversion via Joint Diffusion in the Latent Space

### v2

@@ Abstract
Full Waveform Inversion (FWI) reconstructs high-resolution subsurface velocity maps from seismic waveform data governed by partial differential equations (PDEs). Traditional machine learning approaches frame FWI as an image-to-image translation task, mapping seismic data to velocity maps via encoder-decoder architectures. In this paper, we revisit FWI from a new perspective: generating both modalities simultaneously. We found that both modalities can be jointly generated from a shared latent space using a diffusion process. Remarkably, our jointly generated seismic-velocity pairs inherently satisfy the governing PDE without requiring additional constraints. This reveals an interesting insight: the diffusion process inherently learns a scoring mechanism in the latent space, quantifying the deviation from the governing PDE. Specifically, the generated seismic-velocity pairs with higher scores are closer to the solutions of the governing PDEs. Our experiments on the OpenFWI dataset demonstrate that the generated seismic-velocity pairs not only yield high fidelity, diversity and physical consistency, but also can serve as effective augmentation for training data-driven FWI models.
Full Waveform Inversion, Joint Diffusion, Denoising, Deviation from PDE
@@ 1 Introduction
Subsurface imaging is critical in scientific and industrial applications, including earthquake monitoring (Virieux et al., 2017; Tromp, 2020), greenhouse gas storage (Li et al., 2021; Wang et al., 2023b), medical imaging (Guasch et al., 2020; Lozenski et al., 2024), and oil and gas exploration (Virieux & Operto, 2009; Wang & Alkhalifah, 2018). At its core, subsurface imaging involves reconstructing velocity maps that describe the propagation speed of seismic waves, governed by the acoustic wave equation:
$\frac{\partial^{2}s(\mathbf{x},t)}{\partial t^{2}}=v^{2}(\mathbf{x})\nabla^{2}s(\mathbf{x},t)+f(\mathbf{x_{s}},t),$ (1)
where $s(\mathbf{x},t)$ is the seismic data, $v(\mathbf{x})$ is the velocity model, $\nabla^{2}$ is the Laplacian, and $f(\mathbf{x_{s}},t)$ is the source term. Full Waveform Inversion formulates this task as an inverse problem: given observed seismic data $s(\mathbf{x},t)$, the goal is to estimate the underlying velocity map $v(\mathbf{x})$.
Recently, machine learning-based approaches (Wu & Lin, 2019; Zhang et al., 2019; Sun & Demanet, 2020; Feng et al., 2021) have been introduced to address the limitations (computational cost, cycle skipping, etc) of traditional FWI methods (Plessix, 2006). These machine learning-based methods treat FWI as an image-to-image translation task and use neural networks to directly map from seismic data to velocity maps. In particular, they encode the seismic data $s$ into a latent space $z=E_{\theta}(s)$ from which the velocity map is decoded by $v=D_{\omega}(z)$, where $\theta$ and $\omega$ are the learnable parameters in the encoder and decoder, respectively.
This paper is motivated by an investigation of the latent space $z$. Our preliminary study reveals that randomly sampling from $z$ and computing the corresponding velocity map $v=D(z)$ and seismic reconstruction $\hat{s}=E^{-1}(z)$ (via an additional learned decoder) often results in pairs that do not satisfy the governing PDE in Equation 1. This finding suggests that only a sparse subspace of $z$ corresponds to valid PDE solutions. Naturally, this raises the question: Can we systematically distinguish latent space representations that adhere to the governing PDE from those that do not?
Surprisingly, we found that this can be achieved by applying diffusion models. After training a diffusion network in the shared latent space, we observe that the diffusion process transforms arbitrary latent points into those that correspond to physically valid solutions. This reveals an intriguing insight: the diffusion model automatically learns to score the latent space based on its deviation from the governing PDE. Specifically, higher scores are assigned to latent points from which valid seismic-velocity pairs can be decoded.
Without additional constraints, our framework results in a latent diffusion model that can jointly generate seismic-velocity pairs that approximately satisfy the governing PDE. Experiments on the OpenFWI dataset demonstrate that the generated samples exhibit high fidelity and diversity, providing a valuable source of augmented data for training conventional image-to-image translation models like InversionNet.
The key contribution of this paper is introducing a new generative perspective for FWI: the deviation from the solutions of the governing PDE can be modeled by a diffusion process in the latent space. We believe this insight deepens our understanding of deep learning-based FWI models and bridges generative AI with physics-based modeling.
2 WaveDiffusion: Our Method
@@
In this section, we present WaveDiffusion, a framework designed to explore and refine the latent space extracted by encoder-decoder models for FWI. Specifically, we aim to address two fundamental questions: (1) Do all the points in the latent space constructed by traditional image-to-image mapping models satisfy the governing PDE? (2) If not, can we distinguish latent samples $\{z\}$ that satisfy the governing PDE from those that do not? To investigate these questions, we formulate a two-stage framework: first, we construct the latent space using an encoder-decoder model, then we refine it through a diffusion process to ensure adherence to the governing PDE.
@@ 2.1 Stage 1: Encoder-Decoder and Reconstruction
Achieving comparable performance in FWI: We evaluate the inversion performance of our vector-quantized encoder-decoder model against BigFWI-B (Jin et al., 2024), a state-of-the-art InversionNet trained on the full OpenFWI dataset. This provides a fair comparison, as the training sample volume and model parameter size of BigFWI-B (24M) approximately align with our Stage 1 encoder-decoder model (19M).
Figure 1 shows that our encoder-decoder model achieves competitive performance relative to BigFWI-B across all datasets, outperforming BigFWI-B in six out of ten datasets (marked with yellow stars).
@@
Most latent points do not correspond to PDE solutions: The deviation of decoded modalities from randomly sampled $z$ points that are usually far away from trained latent points, reaches an average L2 distance of $0.013$ (Figure 4 backward step 0), which is eight times higher than the deviation ($0.0016$, Figure 4 forward step 0) of reconstructed modalities from the training set (trained latent points). Figure 2 visually compares the generated samples from Stage 1 (row 1) and the original OpenFWI dataset (row 3). The encoder-decoder provides coarse approximations of PDE solutions but does not fully satisfy the wave equation.
Figure 2: Comparison between generated and original samples: Examples of generated pairs by the (row 1) encoder-decoder and (row 2) joint diffusion. Row 3 shows an original OpenFWI example. The encoder-decoder-generated samples lack the physical relationships governed by the wave equation, while diffusion refines them to better satisfy the PDE.
Thus, encoding and decoding through this latent space alone does not inherently ensure physical validity. Our experiments show that most generated samples fail to satisfy the governing PDE, meaning only a sparse subspace of $z$ corresponds to valid PDE solutions.
@@ 2.2 Stage 2: Joint Diffusion in the Latent Space
Refining the latent space with diffusion: Since random sampling in the latent space does not guarantee adherence to the PDE, we use a joint diffusion model to progressively transform arbitrary latent points into valid ones. This process follows a standard forward-backward diffusion formulation:
@@
Backward process: The noisy latent vector is progressively denoised through the reverse diffusion process: $z_{t-1}=z_{t}-\gamma_{t}\nabla_{z_{t}}\mathcal{L}(z_{t},t)$, where $\gamma_{t}$ is a step size, and $\mathcal{L}$ is the denoising objective defined as:
@@
This follows the standard formulation of denoising score matching for latent diffusion models (Rombach et al., 2022). Here, $\epsilon_{\theta}(z_{t},t)$ is the predicted noise estimated by the learned model. The loss function minimizes the discrepancy between the true noise $\epsilon$ and the model’s predicted noise $\epsilon_{\theta}(z_{t},t)$, guiding the model to iteratively refine the noisy latent variable back to a valid solution in the latent space.
@@ 2.3 Inspecting Diffusion Process
To analyze the role of diffusion, we measure the deviation of generated seismic-velocity pairs from the governing PDE at each diffusion step using the L2 distance between $\hat{s}$ and $s_{\hat{v}}$. Figure 3 visualizes this process, while Figure 4 presents the statistical evaluation of 13,200 generated pairs at each diffusion step.
@@
Figure 3: Visualization of deviation from PDE during diffusion. Seismic data comparison of a CVB example at different timesteps during the forward (left half) and backward diffusion processes (right half). Rows 1-3 show generated seismic channels $\hat{s}$ by the joint diffusion model. Row 4 shows the generated velocity map $\hat{v}$. Rows 5-7 show the ground truth seismic data $s_{\hat{v}}$ calculated for the generated $\hat{v}$. Row 8 shows the deviation from the PDE, visualized as the channel-stacked difference between $s_{\hat{v}}$ and $\hat{s}$. Noise increases the deviation during the forward diffusion, and the reverse process reduces discrepancies.
@@
Deviation increases by adding noise: In the left half of Figure 3, during the forward diffusion process, as noise is added, the seismic data $\hat{s}$ generated by the joint diffusion model diverges more from the ground truth $s_{\hat{v}}$. This divergence, shown as the channel-stacked difference in the last row, reflects the increasing deviation from PDE as the noise level rises. Meanwhile, the decoded modalities show distortion in structures. The statistical evaluation in Figure 4 illustrates how the deviation increases with noise in the forward process, from the initial deviation 0.0016 (trained latent points) to 0.0121 (pure noise points).
@@
Deviation decreases by denoising: In contrast, the right half of Figure 3 shows the backward process, where noise is progressively removed, reducing the deviation and refining the seismic-velocity pairs toward solutions with smaller stacked seismic difference and restored structures. Similarly the statistical evaluation in Figure 4, the deviation decreases as the generated pairs are refined into physically valid solutions from an averaged L2 distance of 0.013 back to 0.002, confirming the model’s ability to refine latent points toward physically consistent solutions.
@@
Figure 4: Deviation from the governing PDE. The L2 distance is calculated between (a) generated seismic data $\hat{s}$ and (b) ground truth seismic data $s_{\hat{v}}$ calculated for the generated velocity map $\hat{v}$ using a finite difference solver.
@@
Our results reveal that only a small subspace of latent points satisfies the PDE, while diffusion refines high-deviation points into physically valid solutions. This suggests that diffusion implicitly scores the latent space, moving it toward PDE-compliant solutions. The diffusion process effectively transforms the deterministic PDE problem into a stochastic differential equation (SDE), allowing exploration of the solution space. This insight provides a new perspective to bridge AI-driven generative modeling with physical principles.
@@ 3 Experiments
In this section, we present the experimental evaluation of the proposed WaveDiffusion framework. We conduct experiments on the full OpenFWI dataset to evaluate the model’s performance in generating physically consistent seismic data and velocity maps. We assess the model’s FID scores and show its ability to improve FWI results. Then, we compare the results of training the state-of-the-art models such as BigFWI using the jointly generated dataset against the original benchmark (Jin et al., 2024). In the Appendix, we further introduce an experiment to demonstrate how the joint diffusion model compares to separately trained diffusion models in generating seismic and velocity modalities.
@@ 3.1 Dataset and Training Setup
In the experiments, we evaluate two different configurations for the latent diffusion models. Specifically, we test (1) two separate VQ codebooks for the two modalities and (2) a single shared VQ codebook for both modalities.
We evaluate the performance of our WaveDiffusion framework using the OpenFWI dataset, a comprehensive benchmark collection comprising 10 subsets of realistic synthetic seismic data paired with subsurface velocity maps, specifically designed for FWI tasks. These subsets represent diverse geological structures, including curved velocity layers, flat velocity layers, and flat layers intersected by faults, among others, allowing for an extensive evaluation of our model’s robustness and generalization capability.
Our experiments utilized all 10 subsets (Fault, Vel, and Style Families) with over 400K training data pairs to ensure a thorough assessment of WaveDiffusion. We performed training on the combined dataset comprising all subsets to assess the model’s ability to generalize across a variety of geological configurations and scenarios. We trained our models on 128 NVIDIA GH200 GPUs. Network details and training hyperparameters used are provided in Appendix A.1.
3.2 Evaluating Generated Samples with FID
The Stage 1 model of our WaveDiffusion framework produces coarse approximations of seismic-velocity pairs. The Stage 1 model without a diffusion process yielded an FID score of 14,207.14 for the randomly sampled $z$ vectors and their corresponding decoded velocity maps and 871.31 for the decoded seismic data using an Inception-v3 model pre-trained on ImageNet (Szegedy et al., 2016). A visualization example is shown in Figure 2 row 1. These high FID scores suggest that the encoder-decoder architecture, while generating plausible shapes, does not adhere closely to the true data distribution. The large disparity between seismic and velocity FID scores indicates that the generated modalities deviate more from the physical relationships governed by the wave equation as coarse approximations of the PDE solutions.
Table 1: FID scores. The FID scores of the two modalities velocity maps $v$ and seismic data $s$ for two model settings.
[…]

### v3

# WaveDiffusion: Joint Latent Diffusion for Physically Consistent Seismic and Velocity Generation

@@ Abstract
Full Waveform Inversion (FWI) is a critical technique in subsurface imaging, aiming to reconstruct high-resolution subsurface properties from surface measurements. Acoustic FWI involves two physical modalities, seismic waveforms and velocity maps, which are governed by the acoustic wave equation. Prior works primarily focus on the inverse problem, modeling the relationship between seismic and velocity as an image-to-image translation task. In this work, we study their relationship from a generative perspective. Our aim is to explore and characterize the latent space structure, and identify latent vectors that generate seismic–velocity pairs consistent with the governing partial differential equation (PDE). Specifically, we model seismic and velocity data jointly from a shared latent space via a diffusion process. In experiments, we find that diffusion progressively refines arbitrary latent vectors into ones that yield approximately physics-consistent seismic–velocity pairs, even without explicit physics constraints. This provides empirical evidence of PDE-consistency in latent diffusion, where sampling is biased toward PDE-valid solutions. In latent space, satisfying the acoustic wave equation can be approximated through sampling and gradient descent. We formalize this physics-consistent latent modeling task and quantify it through extensive experiments. On large-scale OpenFWI benchmarks, our approach produces high-fidelity, diverse, and physically consistent seismic–velocity pairs, demonstrating the potential of a data-driven latent diffusion for physically consistent generation in a complex scientific domain.
Figure 1: Overview of WaveDiffusion. WaveDiffusion generates paired samples that satisfy the governing PDE via a joint latent diffusion framework. The diffusion guides the sampling process toward physically valid solutions through denoising, progressively mapping non-solution points (gray squares in valleys) to valid solutions (colored stars at peaks).
@@ 1 Introduction
Subsurface imaging is critical in scientific and industrial applications, including earthquake monitoring Virieux et al. (2017); Tromp (2020), greenhouse gas storage Li et al. (2021); Wang et al. (2023b), medical imaging Guasch et al. (2020); Lozenski et al. (2024), and oil and gas exploration Virieux and Operto (2009); Wang and Alkhalifah (2018). At its core, Full Waveform Inversion (FWI) involves reconstructing velocity maps that describe the propagation speed of seismic waves, governed by the acoustic wave equation:
$\frac{\partial^{2}s(\mathbf{x},t)}{\partial t^{2}}=v^{2}(\mathbf{x})\nabla^{2}s(\mathbf{x},t)+S(\mathbf{x_{s}},t),$ (1)
where $s(\mathbf{x},t)$ is the seismic data, $v(\mathbf{x})$ is the velocity model, $\nabla^{2}$ is the Laplacian, and $S(\mathbf{x_{s}},t)$ is the source term. Most recent machine learning (ML) approaches have focused on the FWI problem, which formulates this task as an inverse problem. For example, supervised-based methods frame the relationship between seismic and velocity as an image-to-image translation task Wu and Lin (2019); Zhang et al. (2019); Sun and Demanet (2020); Feng et al. (2021).
In this work, we revisit the relationship between seismic and velocity from a generative perspective. While seismic and velocity data are connected by the wave equation, their distribution in the original seismic-velocity joint space is unknown. Rather than treating their relationship purely as an inversion task, our goal is to explore and characterize the latent space structure shared by both modalities. Our preliminary study reveals that randomly sampled latent vectors often decode into pairs that violate the PDE 1, suggesting only a sparse subset of the latent space corresponds to physically valid solutions. Naturally, this raises the question: Can we systematically identify physically valid latent space representations that adhere to the governing PDE from those that do not?
Our key finding is that diffusion models naturally enforce physical consistency when applied in a shared latent space. By jointly modeling seismic data and velocity map from a shared latent space via a diffusion model, the denoising process progressively transforms arbitrary latent sampling points into ones that decode into solutions more consistent with the governing PDE, as illustrated in Figure 1. This indicates that, even without the need for explicit physical constraints, a purely data-driven latent diffusion implicitly biases sampling toward physically valid regions. In other words, in the latent space, approximate satisfaction of the acoustic wave equation can emerge through sampling and gradient descent. Such behavior cannot be assumed a priori and highlights an emergent property of latent diffusion. This data-driven approach is particularly attractive when the forward model is unavailable, non-differentiable, or computationally prohibitive, offering a practical alternative.
Differing from recent diffusion-based approaches to inverse problems (Song et al., 2022; Chung et al., 2023; Zhang et al., 2025), which recover one modality from the other (e.g., velocity from seismic), our objective focuses on the physics-consistent latent modeling task. The goal is to identify latent vectors that jointly generate both modalities in a manner consistent with the governing PDE. More broadly, our work is an exploration into the intersection of generative modeling and scientific data modeling, demonstrating both the potential and the limitations of current approaches.
On large-scale OpenFWI Deng et al. (2022) benchmarks, we evaluate our approach through extensive experiments, demonstrating it produces high-fidelity, diverse, and physically consistent seismic–velocity pairs. Furthermore, we show that these jointly generated pairs can augment training data for conventional supervised models such as InversionNet Wu and Lin (2019), thereby validating both their quality and practical utility.
2 WaveDiffusion: Exploring latent space via joint generation
@@
In this section, we introduce WaveDiffusion, a framework for jointly modeling seismic and velocity data in a shared latent space. Specifically, we first formalize this physics-consistent latent modeling task. Then, we investigate two fundamental questions: (1) Do all latent representations produced by conventional image-to-image models satisfy the governing PDE? (2) If not, can we distinguish latent samples $\{z\}$ that yield physically consistent outputs from those that do not? To address these questions, we adopt a two-stage latent diffusion framework that aligns naturally with our objective. First, we construct a latent space using an encoder-decoder model. Then, we apply a latent diffusion process to refine latent codes toward physically valid solutions progressively.
2.1 Physics-Consistent Latent Modeling
We now formalize the notion of the physics-consistent latent modeling task.
Let $f:\mathcal{V}\to\mathcal{S}$ denote the forward modeling operator (e.g., a numerical wave simulator), mapping a velocity field $v\in\mathcal{V}\subset\mathbb{R}^{H\times W}$ to its corresponding seismic response $s\in\mathcal{S}\subset\mathbb{R}^{T\times W}$, i.e., $s=f(v)$. Let $z\in\mathcal{Z}\subset\mathbb{R}^{d}$ denote a latent representation in a learned space. We assume the existence of two deterministic decoders: $D_{v}:\mathcal{Z}\to\mathcal{V},\quad D_{s}:\mathcal{Z}\to\mathcal{S},$ where $D_{v}$ maps latent vectors to velocity fields and $D_{s}$ maps latent vectors to seismic signals.
A latent vector $z\in\mathcal{Z}$ is defined to be physically valid if it satisfies the following condition:
$\|D_{s}(z)-f(D_{v}(z))\|<\epsilon,$
for some tolerance $\epsilon>0$. In words, the seismic signal decoded from $z$ must be close to the one obtained by applying the forward operator to the decoded velocity. The central challenge is thus to characterize or sample from the subset of the latent space that satisfies this condition:
$\mathcal{Z}_{\text{valid}}:=\{z\in\mathcal{Z}\mid\|D_{s}(z)-f(D_{v}(z))\|<\epsilon\}.$
Two fundamental properties of this problem are of particular interest:
  • Completeness. To what extent can the valid set $\mathcal{Z}_{\text{valid}}$ be captured? That is, what fraction of physically valid solutions are representable under the learned distribution $p(z)$?
  • PDE Satisfaction. How small can the residual error $\epsilon$ be made in practice, and what factors (e.g., model capacity, dataset coverage, optimization dynamics) govern this error bound?
In practice, as we will discuss in the following, latent vectors sampled arbitrarily from $\mathcal{Z}$ typically fail to meet this criterion, which motivates the need to learn a structured distribution $p(z)$ over $\mathcal{Z}_{\text{valid}}$.
@@ 2.2 Stage 1: Encoder-Decoder and Reconstruction
Achieving comparable performance in FWI: We evaluate the inversion performance of our vector-quantized encoder-decoder model against BigFWI-B Jin et al. (2024), a state-of-the-art InversionNet trained on the full OpenFWI dataset. This provides a fair comparison, as the training sample volume and model parameter size of BigFWI-B (24M) approximately align with our Stage 1 model (19M). Figure 2 shows that our encoder-decoder model achieves competitive performance relative to BigFWI-B across all datasets, outperforming BigFWI-B in six out of ten datasets (marked with yellow stars).
@@
Most latent points do not correspond to PDE solutions: The deviation of decoded modalities from randomly sampled $z$ points, which are usually far away from trained latent points, reaches an average L2 distance of $0.013$ (Figure 5 backward step 0), which is eight times higher than the deviation ($0.0016$, Figure 5 forward step 0) of reconstructed modalities from the training set (trained latent points). Figure 3 visually compares the generated samples from Stage 1 (row 1) and the original OpenFWI dataset (row 3). The encoder-decoder provides coarse approximations of PDE solutions but does not fully satisfy the wave equation.
Figure 3: Generated samples Comparison.
To address whether poor reconstructions result merely from latent distribution mismatch, we conduct additional experiments, detailed in Appendix A.3, and further confirm that the diffusion model significantly improves reconstruction quality. Encoding and decoding through this latent space alone do not inherently ensure physical validity. Thus, most generated samples fail to satisfy the governing PDE, meaning only a sparse subset of $z$ corresponds to valid solutions.
@@ 2.3 Stage 2: Joint Diffusion in the Latent Space
Refining the latent space with diffusion: Since random sampling in the latent space does not guarantee adherence to the PDE, we use a joint diffusion model to transform arbitrary latent points into valid ones progressively. This process follows a standard forward-backward diffusion formulation:
@@
Backward process: The noisy latent vector is progressively denoised through the reverse diffusion process: $z_{t-1}=z_{t}-\gamma_{t}\nabla_{z_{t}}\mathcal{L}(z_{t},t)$, where $\gamma_{t}$ is a step size, and $\mathcal{L}$ is the denoising objective:
@@
This follows the standard formulation of denoising score matching for latent diffusion models Rombach et al. (2022). Here, $\epsilon_{\theta}(z_{t},t)$ is the predicted noise estimation. The loss function minimizes the discrepancy between the true noise $\epsilon$ and the model’s predicted noise $\epsilon_{\theta}(z_{t},t)$, guiding the model to iteratively refine the noisy latent variable back to a valid solution in the latent space.
@@ 2.4 Inspecting Diffusion Process
To analyze the role of diffusion, we measure the deviation of generated seismic-velocity pairs from the governing PDE at each diffusion step using the L2 distance between $\hat{s}$ and $s_{\hat{v}}$. Figure 4 visualizes this process, while Figure 5 presents the statistical evaluation on 13.2k samples at each diffusion step.
@@
Figure 4: Visualization of deviation from PDE during diffusion. Seismic data of a CVB example at different timesteps during the forward (left half) and backward diffusion processes (right half). Rows 1-3 show generated seismic channels $\hat{s}$ by the joint diffusion model. Row 4 shows the generated velocity map $\hat{v}$. Rows 5-7 show the ground truth seismic data $s_{\hat{v}}$ calculated for the generated $\hat{v}$. Row 8 shows the deviation, visualized as the channel-stacked difference between $s_{\hat{v}}$ and $\hat{s}$. Noise increases the deviation during the forward diffusion, and the reverse process reduces discrepancies.
@@
Deviation increases by adding noise: In the left half of Figure 4, during the forward diffusion process, as noise is added, the seismic data $\hat{s}$ generated by the model diverges more from the ground truth $s_{\hat{v}}$. This divergence, shown as the channel-stacked difference in the last row, reflects the increasing deviation from PDE as the noise level rises. Meanwhile, the decoded modalities show distortion in structures. The statistical evaluation in Figure 5 illustrates how the deviation increases, from the initial deviation 0.0016 (trained latent points) to 0.0121 (pure noise points).
@@
Deviation decreases by denoising: In contrast, the right half of Figure 4 shows the backward process, where noise is progressively removed, reducing the deviation and refining the seismic-velocity pairs toward solutions with smaller stacked seismic difference and restored structures. Similarly, the statistical evaluation in Figure 5, the deviation decreases as the generated pairs are refined into physically valid solutions from an averaged L2 distance of 0.013 back to 0.002, confirming the model’s ability to refine latent points toward physically consistent solutions.
@@
Figure 5: Deviation from the governing PDE. The L2 distance is calculated between generated seismic data $\hat{s}$ and ground truth $s_{\hat{v}}$ of the generated velocity map $\hat{v}$ using a finite difference solver.
@@
Our results reveal that only a small subset of latent points satisfies the PDE, while diffusion progressively moves high-deviation points toward physically valid solutions. This suggests that diffusion implicitly evaluates the latent space, guiding toward PDE-compliant solutions. Moreover, the diffusion process can be viewed as transforming the deterministic PDE constraint into a stochastic differential equation (SDE), thereby enabling exploration of the solution space and providing a new perspective for bridging generative modeling with physical principles.
Building on this insight, we can empirically examine the performance of WaveDiffusion with respect to the two central properties of physics-consistent latent modeling. Completeness. Diffusion models sample from the denoising distribution learned during training; thus, their coverage is constrained by the support of the training distribution and the noise schedule. PDE Satisfaction. We find that diffusion provides a close approximation to the underlying distribution $p(z)$. The dominant residual error arises from the VQGAN decoder rather than the diffusion model itself, as illustrated by the alignment observed at the beginning and end of the diffusion process in Figure 5.
@@ 3 Experiments
In this section, we present the experimental evaluation of the proposed WaveDiffusion framework. We conduct experiments on the full OpenFWI dataset to evaluate the model’s performance in generating physically consistent seismic data and velocity maps. We assess the model’s FID scores and show its ability to improve FWI results. Then, we compare the results of training the state-of-the-art models such as BigFWI using the jointly generated dataset against the original benchmark Jin et al. (2024). In the Appendix, we further introduce an experiment to demonstrate how the joint diffusion model compares to separately trained diffusion models in generating both two modalities.
@@ 3.1 Dataset and Training Setup
In the experiments, we evaluate two different configurations for the latent diffusion models: (1) two separate VQ codebooks for the two modalities and (2) a single shared VQ codebook for both modalities. We evaluate the performance of our WaveDiffusion framework using the OpenFWI dataset, a comprehensive benchmark collection comprising 10 subsets of realistic synthetic seismic data paired with subsurface velocity maps, specifically designed for FWI tasks. These subsets represent diverse geological structures, including curved velocity layers, flat velocity layers, and flat layers intersected by faults, among others, allowing for an extensive evaluation of our model’s robustness and generalization capability.
Our experiments utilized all 10 subsets (Fault, Vel, and Style Families) with over 400K training data pairs to ensure a thorough assessment of WaveDiffusion. We performed training on the combined dataset comprising all subsets to assess the model’s ability to generalize across a variety of geological configurations and scenarios. We trained our models on 16 NVIDIA H100 GPUs. Network details and training hyperparameters used are provided in Appendix A.1.
Table 1: FID scores. The FID scores of velocity maps $v$ and seismic data $s$ for two model settings.
Metrics $\backslash$ Model 2VQ 1VQ
@@
3.2 Evaluating Generated Samples with FID
The Stage 1 model of our WaveDiffusion framework produces coarse approximations of seismic-velocity pairs. The Stage 1 model without a diffusion process yielded an FID score of 14,207.14 for the randomly sampled $z$ vectors and their corresponding decoded velocity maps and 871.31 for the decoded seismic data using an Inception-v3 model pre-trained on ImageNet Szegedy et al. (2016). A visualization example is shown in Figure 3 row 1. These high FID scores suggest that the encoder-decoder architecture, while generating plausible shapes, does not adhere closely to the true data distribution. The large disparity between seismic and velocity FID scores indicates that the generated modalities deviate more from the physical relationships governed by the wave equation as coarse approximations of the PDE solutions.
Figure 6: t-SNE visualization for both real (blue) and generated (orange) data. Results of seismic data (col 1) and velocity map (col 2) for two joint diffusion models.